OpenStudy (anonymous):

Katrina is building a deck. She cuts 3 lengths of wood from a 72-inch board. The first length is 30 3/4 inches. The second length is 12 1/4 inches. And the third length is 16 1/4 inches. How much of the original board remains after Katrina makes her cuts?

OpenStudy (anonymous):

\[72-\left(30+\frac{3}{4}\right)-\left(12+\frac{1}{4}\right)-\left(16-\frac{1}{4}\right)=\frac{53}{4}=13+\frac{1}{4} \text{ inches}\]